OpenStudy (anonymous):

Sry again :/ Physics. Medal + Fan Thanks :') A drum vibrates 180 times in 2 seconds. If the speed of sound is 340m/s, what is the period, frequency and wavelength of the waves produced by the drum? The speed of light is 3 X 10^8m/s.

OpenStudy (loser66):

as you said in the previous problem, frequency = 180/2 =90

OpenStudy (loser66):

period= 1/f = 1/90

OpenStudy (anonymous):

= 0.01

OpenStudy (loser66):

and the wavelength is calculated by \[v=\dfrac{\lambda}{period} \\\lambda = v *period\] you force me look up my physics book which I "swear" I never touch after I passed the course. hehehe @dan815 he is engineer expert

OpenStudy (anonymous):

:D tanks :;)